OpenStudy (blowery):

Convert 5.6cm^3 to mL help??

OpenStudy (yanasidlinskiy):

@agent0smith is good at this... Wouldn't it be 5.6?

OpenStudy (astrophysics):

1 cubic cm = 1 ml

OpenStudy (blowery):

Is that because volume is implied to be cubed already?

OpenStudy (astrophysics):

So yes @YanaSidlinskiy is correct

OpenStudy (astrophysics):

Yeah you can think of volume as area*height that's what it really is

OpenStudy (astrophysics):

area is 2 dimensional multiply by height you get 3d

OpenStudy (blowery):

You guys are awesome. What about 6.7 kL to cm^3. Would that be 6,700,000

OpenStudy (astrophysics):

Yeah that looks good, units are very important so a way you can do it is \[\left( 6.7 kL \right) \times \left( \frac{ cm^3 }{ 10^{-6} kL } \right) = 6.7 \times 10^6 cm^3\]

OpenStudy (astrophysics):

Notice the kL gets cancelled out

OpenStudy (blowery):

We really haven't gotten to that formula, but I'm certain we will. I just replaced 1 cm^3 with mL and then moved the decimal six places to the right, 3 places to get from kL to L and three more to get from L to mL. Is my reasoning sound?

OpenStudy (astrophysics):

It's not a formula I just knew that in 1 cm^3 there is 10^-6 kL and then you can easily cancel out the units, yours seems fine but a lot of work but if it works for you heck go for it!

OpenStudy (astrophysics):

If you google dimensional analysis you will find a lot of info on how to do these problems

OpenStudy (blowery):

Haha, I was thinking yours seemed like a lot of work. But I really appreciate your help and thank you for showing me a better way!

OpenStudy (astrophysics):

No problem, as to my previous post dealing with units is called "dimensional analysis" and you need it everywhere :)